OpenStudy (anonymous):

If the two legs of a right triangle are 5 and 4, what is the length of the hypotenuse? A. 3 B. 9 C. 41 D. √41

OpenStudy (oaktree):

5^2+4^2=25+16=41, but square root is sqrt(41).

OpenStudy (oaktree):

So D.